OpenStudy (wwwhhhaaattt?):

Just for fun. I will metal who can answer all of these correct You have to do it within 5 min My grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years ? 2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Using eight eights and addition only, can you make 1000?

OpenStudy (studygurl14):

Again, the last one is the easiest. 888+8+88+8+8=1000

OpenStudy (wwwhhhaaattt?):

DAMIT WHY DO YOU HAVE TO BE SO SMART

OpenStudy (queelius):

Solve the system of equations: \[x_1=(1/7)x_2\] \[(1/30)x_1=(1/360)x_3\] \[(x_1+x_2+x_3)/365=120\] Not accounting for leap years, and things like that, since we don't know the date of births for any of them. Note that x1 is grandson's age in days, x2 is son's age in days, and x3 is your age in days.

OpenStudy (wwwhhhaaattt?):

Ok

OpenStudy (queelius):

I'm not going to bother to solve it. It's a system of equation; 3 unknowns, and 3 equations, so it can be solved (assuming there is a solution).

OpenStudy (wwwhhhaaattt?):

Well ill give it to you anyway for trying.

OpenStudy (queelius):

Err, 1/360 in equation two should be 1/365

OpenStudy (queelius):

It's a straightforward calculation (simple substitution), it's just tedious. Not interested. :)